OpenStudy (anonymous):

An object is traveling around a circle with a radius of 10 cm. If in 20 seconds a central angle of 1/3 radian is swept out, what is the linear speed of the object? I think I've set my formula up right (v=Δ20/Δ1/3 radian), but I don't know what to do with the Δ thing. Once I know how to do that I should be able to like, actually figure it out.

OpenStudy (nurali):

ang_speed = (1/3 rad) / (20 s) = 1/60 rad/s lin_speed = R*(ang_speed) = 10 * (1/60) = 1/6 cm/s

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

Ahh, wow thanks! I kind of feel like an idiot because it was so simple.

OpenStudy (nurali):

My Pleasure.